# Superspin Renormalization and Slow Relaxation in Random Spin Systems

### v1

@@ I Introduction
Recent years have witnessed the development of experimental platforms enabling high-fidelity local control over and readout from many-body spin systems. Many of these systems, such as nitrogen-doped diamond [1, 2, 3, 4] as well as arrays of Rydberg atoms [5], ultracold magnetic atoms [6, 7], and polar molecules [8], feature long-range, often dipolar interactions between spins that are randomly or arbitrarily positioned. These developments call for progress in theories that can yield predictions for the dynamics of systems with irregular geometries and long-range interactions.
@@
In this work, we study the late-time local relaxation of conserved densities in such systems. For concreteness, we examine the relaxation of the conserved spin density in systems with random $XX+YY$ interactions. We discover a dramatic impact of the randomness on the infinite-temperature autocorrelation function of the $Z$ component of the spin, whose average over realizations of the randomness (or, in a large system, over space) features heavy late-time tails decaying slower than any power law. In experiments, this autocorrelation function can be measured after the preparation of randomly initialized spins and has served as a probe of their ensuing dynamics [4, 3, 9]. We put an emphasis on chains with dipolar $XX+YY$ interactions and positional randomness, as are particularly relevant to experiments on Rydberg atoms [5], including for example early work emulating a three-site spin chain [10].
@@
A longer version of such an array of Rydberg atoms, with some fraction of sites left empty as typically occurs in experiments, will turn out to realize a kind of random quantum Hamiltonian that has been of great theoretical interest for decades. The analysis of disordered Hamiltonians has been generally advanced by real-space, strong-randomness renormalization group (RSRG) approaches [11, 12, 13, 14]. Particularly, the ground state and low-energy excited states of one-dimensional spin chains under random nearest-neighbor interactions have been intensively studied with RSRG and shown to possess universal properties in many cases [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. The above-mentioned experiments, however, can hardly access low temperatures due to the lack of clear cooling protocols.
@@
On the theory side, several works have generalized the ground-state RSRG to describe, under specific setups, generic excited states (RSRG-X) or non-equilibrium dynamics [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. An RSRG-X formalism centers on an iterative scheme to update an effective Hamiltonian $H_{\mathrm{eff}}$ to lower energy scales. In a random system, these updates involve identifying a strong coupling in $H_{\mathrm{eff}}$ and approximating all others as perturbations. One then updates $H_{\mathrm{eff}}$ to describe dynamics within each eigenspace of the strong coupling. Physically, such treatment can be viewed as “integrating out” the fast degrees of freedom. Typically, the randomness of the system improves along RSRG-X, leading to more trustworthy predictions at lower energy scales and hence later times. We propose an RSRG-X formalism to analyze infinite-temperature spin relaxation under the experimentally relevant setups.
@@
Our work extends previous efforts to study the dynamics of randomly interacting spin chains with RSRG approaches. Early works have focused on low-temperature dynamics [21, 23, 24, 26], where it has been argued that nontrivial superspins do not play a significant role [24]. The non-equilibrium quench dynamics of interacting spins chains respecting $\text{U}(1)$ and $\mathbb{Z}_{2}$ symmetries have been later studied using RSRG-X in Ref. [29]; there, the initial states are chosen to suppress contributions from nontrivial superspins. RSRG-X techniques have since been developed for Hamiltonians that are sums of Pauli strings (i.e., tensor products of Pauli matrices $X,Y,Z$) that have statistically independent strengths [30, 31, 34, 36], as well as for nearest-neighbor Heisenberg models [37]. Unfortunately, these RSRG-X developments do not apply to the general (including long-range) random $XX+YY$ models of our interest. For nearest-neighbor $XX+YY$ models, Ref. [32] has established an RSRG-X scheme for all of the excited states and calculated their entanglement entropies, while in models which do not map to non-interacting fermions Refs. [38, 39, 40, 41] have formulated simplified RSRG-X schemes which neglect contributions from nontrivial superspins.
@@ II.1 Models
To guide our investigation of the dipolar chain, we first study its nearest-neighbor counterparts, i.e., models with $J_{jk}=0$ for $|j-k|\geq 2$. Here, we label the sites so that $0\leq x_{1}<x_{2}<\cdots<x_{N}<L$, the indices being defined modulo $N$ in light of PBC. Such models admit a simplified RSRG-X formalism leading to analytically controlled analyses, a simplification also expected from the free-fermion nature. There are two types of randomness that we consider, (i) $J_{j,j+1}$ drawn from the “fixed point” distribution, whose functional form is invariant under the analytic RSRG-X (see Equ. (19) and Sec. IV.1); and (ii) $J_{j,j+1}$ as in the dipolar model. Additionally, we study the next-nearest-neighbor model where interactions $J_{jk}$ with $|j-k|\geq 3$ are truncated from the dipolar chain.
@@ II.2 Autocorrelation Functions
We calculate $\overline{\mathcal{S}_{p}}(\omega)$ using RSRG-X for nearest-neighbor chains in Secs. IV.1 and IV.2, and for next-nearest-neighbor and long-range chains in Sec. V.1. In all cases, we observe low-frequency behaviors that are consistent with the one in Equ. (7). Our RSRG-X formalism is not expected to capture infinite-time behaviors, and hence we should not expect an accurate estimate for the baseline $B$. However, we find remarkable quantitative agreement between the RSRG-X and ED results at late times up to a baseline shift, or equivalently in the regime of low but nonzero frequencies. This agreement holds even when there are significant finite-size effects in both results. In Sec. VI, we extend the analysis to two-dimensional long-range models: we find such agreement only for rapidly decaying interactions, and our results at small system sizes suggest subdiffusive spin relaxation.
@@ III.2.1 The RG Steps
where $J_{LR}=J_{L}J_{R}/\Omega$. Note that, for all cases, the updated $H_{\mathrm{eff}}$ does not hybridize $\mathcal{H}_{\overline{12}}\otimes\ket{\uparrow_{1}\uparrow_{2}}$ with $\mathcal{H}_{\overline{12}}\otimes\ket{\downarrow_{1}\downarrow_{2}}$ and still takes the form of a nearest-neighbor $XX+YY$ chain. Consequently, we can view the updated $H_{\mathrm{eff}}$ as consisting of two fewer sites — the sites $1$, $2$ are cast out, and we fix the degrees of freedom on these two sites to be one of the four eigenstates of $H_{0}$ (particularly for branching into $V_{0}$, we choose between |$\ket{\uparrow_{i}\uparrow_{j}}$, |$\ket{\downarrow_{i}\downarrow_{j}}$ with definite magnetization (2), although in principle a linear combination would also work as they do not hybridize (13)). A new interaction of strength $|J_{LR}|$ is generated between the sites $L$ and $R$, which now become nearest neighbors. One can then start another RG step with the updated $H_{\mathrm{eff}}$ and repeat the procedures above.
@@ III.2.3 An Analytical Description
The flow of $P_{\Gamma}(|J|)$ at late RG steps has been first derived in the ground-state RSRG for the antiferromagnetic chain (i.e., Equ. (10)) with $J_{j}>0$) [19]: at every RG step, branching into the ground-state subspace $V_{-}$ always introduces a weaker antiferromagnetic interaction $J_{LR}=J_{L}J_{R}/\Omega$ between the neighboring sites (13). It is convenient to introduce the logarithmic interaction strengths $\beta_{j}=\log(|\Omega/J_{j}|)$ so that the update rule reads $\beta_{LR}=\beta_{L}+\beta_{R}$ and that at every step the strongest interaction $\Omega$ corresponds to $\beta=0$. We can then write down an integro-differential equation describing the flow of $P_{\Gamma}(\beta)$,
@@ IV.2.2 Truncated Dipolar Interactions
To highlight, for the nearest-neighbor $XX+YY$ chain (10) with the two types of randomness realizations (Secs. IV.2.1, Sec. IV.2.2), we observe quantitative agreement (Figs. 2, 4) between numerical RSRG-X and ED results for the low-frequency spin relaxation dynamics at all system sizes. The agreement holds even at small $N$, indicating that RSRG-X has the potential as an efficient numerical approach for predicting late-time dynamics. We will see such power of RSRG-X again in the interacting models (Sec. V), where ED calculations are limited to very small systems.
@@ VI The Two-Dimensional Model
Additionally, both the ED and RSRG-X results in Fig. 12 suggest power-law decay of $\overline{S_{p}}(t)\sim t^{-c}$ at late times. The extracted exponent $c$ indicates slower decay at larger $\alpha$, and for all $\alpha\geq 3$ we observe subdiffusive relaxation; note that diffusion would correspond to $c=1$ in two dimensions. We emphasize that, admittedly, the system size we have access to is too small to lead to definite conclusions in the thermodynamic limit.
@@ VII Conclusion
Our work raises general questions on phenomenology of spin relaxation in one- and higher-dimensional systems with random dipolar interactions. In one dimension, our results for the next-nearest-neighbor model (1) suggest no deviation from the late-time $\overline{S_{p}}(t)-\overline{S_{p}}(\infty)\sim\left[\log(t)\right]^{-2}$ asymptote established for the nearest-neighbor model. For the long-range model, our RSRG-X simulation suffers from strong finite-size effects, but we observe the same asymptotic behavior at the largest accessible system size $N=40$ (Fig. 10). In both cases, our results suggest slower-than-power-law decay up to surprisingly large timescale in one dimension; recall that the typical energy scale (1) is fixed to unity. In two dimensions, we observe power-law, sub-diffusive decay of $\overline{S_{p}}(t)$ in the small system sizes to which we have access (Fig. 12). Further numerical and even analytical advances are necessary to determine the phenomenology in the thermodynamic limit.
@@
The presence of slow dynamics due to randomness has been explored in numerous works, including recent modifications of Lieb-Robinson bounds [52, 53] and proofs of sub-diffusive transport [54], both concerning disordered spin chains with general types of local interactions. One question to ask is whether the slower-than-power-law decay of spin survival probability we have observed in a family of spin chains with $\mathrm{U}(1)$ and $\mathbb{Z}_{2}$ symmetries suggests possibly tighter bounds on transport properties in this setting. Such constraints on transport, if any, might be advanced by our RSRG-X formalism combined with a careful analysis of the subleading contributions from the perturbative treatments; indeed the general idea that strong interactions renormalize in the first step to effectively weaker ones has been applied in Ref. [53]. Moreover, although Lieb-Robinson bounds have been derived for generic systems with long-range interactions [55, 56], our work raises questions over how these bounds are modified in systems with positional randomness.
@@
On a practical level, the wide range of interacting models to which our RSRG-X applies include systems with long-range random interactions of the form $XX+YY$ and also even $ZZ$. Many cutting-edge experimental platforms feature such interactions and thus can probe the signatures of spin dynamics predicted by our formalism. Conversely, an important application of our method is to calibrate and benchmark emerging hardware for quantum computation. Especially, the quantum simulation of random one-dimensional spin models will be a valuable step forward: the dynamics in these models involve interference effects of a kind not seen in translation-invariant settings. Meanwhile, random systems in one dimension are typically not as intractable as in higher dimensions; even finding the ground state of a random classical Ising model is NP-complete in two dimensions [57]. The rapid progress in experimental capabilities offers the promise that theory and experiment on random quantum spin models can be more strongly coupled than they have been in the past, leading to new discoveries in this rich area of physics.
@@ Appendix B Interactions at two superspin-0 sites
In this appendix, we explain why we can expect the $X^{(0)}X^{(0)}$ and $Y^{(0)}Y^{(0)}$ types of interactions between two superspin-0 sites to be rarely the strongest in the effective Hamiltoinan $H_{\mathrm{eff}}$ updated along RSRG-X. Our reasoning consists of two parts. First, we point out that the superspin-0 sites themselves are rare, so the strongest interaction rarely involves two such sites. Furthermore, we argue that the $X^{(0)}X^{(0)}$ and $Y^{(0)}Y^{(0)}$ interactions are weak if the three-site interactions, as we also expect, are weak. These theoretical considerations are further supported by the numerical RSRG-X of the specific models (1) we consider (Appendix E).
@@
To argue that the $X^{(0)}X^{(0)}$ and $Y^{(0)}Y^{(0)}$ interactions are weak as long as the three-site interactions in $H_{\mathrm{eff}}$ are weak, we consider how such interactions can be generated in some prior RG step. In fact, a superspin-0 site itself can only be generated when we branch into the $V_{-}^{\prime}$ subspace (case (ii) in Sec. V.1.3): let the strongest interaction $H_{0}$ in $H_{\mathrm{eff}}$ at this step involve, without loss of generality, the sites $1$, $2$ of equal superspin-$\frac{m}{2}$ with $m>0$. Then, $H_{\mathrm{eff}}^{(V_{-}^{\prime})}$ contains in replacement of the sites 1, 2 a new superspin-0 site, which we label by 00, as well as new interactions of the form $X^{(0)}X^{(0)}$ and $Y^{(0)}Y^{(0)}$ — let us consider a new interaction, for example, $X_{0}^{(0)}X_{3}^{(0)}$. In light of the $\mathrm{U}(1)$ and $\mathbb{Z}_{2}$ symmetries (31), we exhaust the allowed interactions in $H_{\mathrm{eff}}$ among the sites 1, 2, 3,

### v2

@@ I Introduction
Recent years have witnessed the development of experimental platforms enabling high-fidelity local control over and readout from many-body spin systems. Many of these systems, such as nitrogen-doped diamond [1, 2, 3, 4] as well as arrays of Rydberg atoms [5], ultracold magnetic atoms [6, 7, 8], and polar molecules [9], feature long-range, often dipolar interactions between spins that are randomly or arbitrarily positioned. These developments call for progress in theories that can yield predictions for the dynamics of systems with irregular geometries and long-range interactions.
@@
In this work, we study the late-time local relaxation of conserved densities in such systems. For concreteness, we examine the relaxation of the conserved spin density in systems with random $XX+YY$ interactions. We discover a dramatic impact of the randomness on the infinite-temperature autocorrelation function of the $Z$ component of the spin, whose average over realizations of the randomness (or, in a large system, over space) features heavy late-time tails decaying slower than any power law. In experiments, this autocorrelation function can be measured after the preparation of randomly initialized spins and has served as a probe of their ensuing dynamics [4, 3, 10, 11]. We put an emphasis on chains with dipolar $XX+YY$ interactions and positional randomness, as are particularly relevant to experiments on Rydberg atoms [5], including for example early work emulating a three-site spin chain [12].
@@
A longer version of such an array of Rydberg atoms, with some fraction of sites left empty as typically occurs in experiments, will turn out to realize a kind of random quantum Hamiltonian that has been of great theoretical interest for decades. The analysis of disordered Hamiltonians has been generally advanced by real-space, strong-randomness renormalization group (RSRG) approaches [13, 14, 15, 16]. Particularly, the ground state and low-energy excited states of one-dimensional spin chains under random nearest-neighbor interactions have been intensively studied with RSRG and shown to possess universal properties in many cases [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. The above-mentioned experiments, however, can hardly access low temperatures due to the lack of clear cooling protocols.
@@
On the theory side, several works have generalized the ground-state RSRG to describe, under specific setups, generic excited states (RSRG-X) or non-equilibrium dynamics [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. An RSRG-X formalism centers on an iterative scheme to update an effective Hamiltonian $H_{\mathrm{eff}}$ to lower energy scales. In a random system, these updates involve identifying a strong coupling in $H_{\mathrm{eff}}$ and approximating all others as perturbations. One then updates $H_{\mathrm{eff}}$ to describe dynamics within each eigenspace of the strong coupling. Physically, such treatment can be viewed as “integrating out” the fast degrees of freedom. Typically, the randomness of the system improves along RSRG-X, leading to more trustworthy predictions at lower energy scales and hence later times. We propose an RSRG-X formalism to analyze infinite-temperature spin relaxation under the experimentally relevant setups.
@@
Our work extends previous efforts to study the dynamics of randomly interacting spin chains with RSRG approaches. Early works have focused on low-temperature dynamics [23, 25, 26, 28], where it has been argued that nontrivial superspins do not play a significant role [26]. The non-equilibrium quench dynamics of interacting spins chains respecting $\text{U}(1)$ and $\mathbb{Z}_{2}$ symmetries have been later studied using RSRG-X in Ref. [32]; there, the initial states are chosen to suppress contributions from nontrivial superspins. RSRG-X techniques have since been developed for Hamiltonians that are sums of Pauli strings (i.e., tensor products of Pauli matrices $X,Y,Z$) that have statistically independent strengths [33, 34, 37, 39], as well as for nearest-neighbor Heisenberg models [40]. Unfortunately, these RSRG-X developments do not apply to the general (including long-range) random $XX+YY$ models of our interest. For nearest-neighbor $XX+YY$ models, Ref. [35] has established an RSRG-X scheme for all of the excited states and calculated their entanglement entropies, while in models which do not map to non-interacting fermions Refs. [30, 41, 42, 43, 44, 45] have formulated simplified RSRG-X schemes which neglect contributions from nontrivial superspins.
@@ II.1 Models
To guide our investigation of the dipolar chain, we first study its nearest-neighbor counterparts, i.e., models with $J_{jk}=0$ for $|j-k|\geq 2$. Here, we label the sites so that $0\leq x_{1}<x_{2}<\cdots<x_{N}<L$, the indices being defined modulo $N$ in light of PBC. Such models admit a simplified RSRG-X formalism leading to analytically controlled analyses, a simplification also expected from the free-fermion nature. There are two types of randomness that we consider, (i) $J_{j,j+1}$ drawn from the “fixed point” distribution, whose functional form is invariant under the analytic RSRG-X [see Equ. (19) and Sec. IV.1]; and (ii) $J_{j,j+1}$ as in the dipolar model. Additionally, we study the next-nearest-neighbor model where interactions $J_{jk}$ with $|j-k|\geq 3$ are truncated from the dipolar chain.
@@ II.2 Autocorrelation Functions
We calculate $\overline{\mathcal{S}_{p}}(\omega)$ using RSRG-X for nearest-neighbor chains in Secs. IV.1IV.2, and for next-nearest-neighbor and long-range chains in Sec. V.1. In all cases, we observe low-frequency behaviors that are consistent with the one in Equ. (7). Our RSRG-X formalism is not expected to capture infinite-time behaviors, and hence we should not expect an accurate estimate for the baseline $B$. However, we find remarkable quantitative agreement between the RSRG-X and ED results at late times up to a baseline shift, or equivalently in the regime of low but nonzero frequencies. This agreement holds even when there are significant finite-size effects in both results. In Sec. VI, we extend the analysis to two-dimensional long-range models: we find such agreement only for rapidly decaying interactions, and our results at small system sizes suggest subdiffusive spin relaxation.
@@ III.2.1 The RG Steps
where $J_{LR}=J_{L}J_{R}/\Omega$. Note that, for all cases, the updated $H_{\mathrm{eff}}$ does not hybridize $\mathcal{H}_{\overline{12}}\otimes\ket{\uparrow_{1}\uparrow_{2}}$ with $\mathcal{H}_{\overline{12}}\otimes\ket{\downarrow_{1}\downarrow_{2}}$ and still takes the form of a nearest-neighbor $XX+YY$ chain. Consequently, we can view the updated $H_{\mathrm{eff}}$ as consisting of two fewer sites — the sites $1$, $2$ are cast out, and we fix the degrees of freedom on these two sites to be one of the four eigenstates of $H_{0}$. [Particularly for branching into $V_{0}$, we choose between |$\ket{\uparrow_{i}\uparrow_{j}}$, |$\ket{\downarrow_{i}\downarrow_{j}}$ with definite magnetization (2), although in principle a linear combination would also work as they do not hybridize (13).] A new interaction of strength $|J_{LR}|$ is generated between the sites $L$ and $R$, which now become nearest neighbors. One can then start another RG step with the updated $H_{\mathrm{eff}}$ and repeat the procedures above.
@@ III.2.3 An Analytical Description
The flow of $P_{\Gamma}(|J|)$ at late RG steps has been first derived in the ground-state RSRG for the antiferromagnetic chain [i.e., Equ. (10) with $J_{j}>0$[21]: at every RG step, branching into the ground-state subspace $V_{-}$ always introduces a weaker antiferromagnetic interaction $J_{LR}=J_{L}J_{R}/\Omega$ between the neighboring sites (13). It is convenient to introduce the logarithmic interaction strengths $\beta_{j}=\log(|\Omega/J_{j}|)$ so that the update rule reads $\beta_{LR}=\beta_{L}+\beta_{R}$ and that at every step the strongest interaction $\Omega$ corresponds to $\beta=0$. We can then write down an integro-differential equation describing the flow of $P_{\Gamma}(\beta)$,
@@ IV.2.2 Truncated Dipolar Interactions
To highlight, for the nearest-neighbor $XX+YY$ chain (10) with the two types of randomness realizations (Secs. IV.2.1IV.2.2), we observe quantitative agreement (Figs. 2, 4) between numerical RSRG-X and ED results for the low-frequency spin relaxation dynamics at all system sizes. The agreement holds even at small $N$, indicating that RSRG-X has the potential as an efficient numerical approach for predicting late-time dynamics. We will see such power of RSRG-X again in the interacting models (Sec. V), where ED calculations are limited to very small systems.
@@ VI The Two-Dimensional Model
Additionally, both the ED and RSRG-X results in Fig. 12 suggest power-law decay of $\overline{S_{p}}(t)\sim t^{-c}$ at late times. The extracted exponent $c$ indicates slower decay at larger $\alpha$, and for all $\alpha\geq 3$ we observe subdiffusive relaxation; note that diffusion would correspond to $c=1$ in two dimensions. We emphasize that, admittedly, the system size we have access to is too small to lead to definite conclusions in the thermodynamic limit. One possible way to lift such uncertainty is through experimental studies: a recent experiment on Rydberg atoms randomly positioned in three dimensions has studied the relaxation of the transverse (rather than conserved) component of the spin degrees of freedom and observed stretched-exponential decay of the corresponding autocorrelation function [11]. Our numerical results here in two dimensions, however, suggest qualitatively different behavior for the conserved component. To what extent the spatial dimensions might result in distinct regimes of late-time relaxation dynamics is worth further investigation.
@@ VII Conclusion
Our work raises general questions on the phenomenology of spin relaxation in one- and higher-dimensional systems with random dipolar interactions. In one dimension, our results for the next-nearest-neighbor model (1) suggest no deviation from the late-time $\overline{S_{p}}(t)-\overline{S_{p}}(\infty)\sim\left[\log(t)\right]^{-2}$ asymptote established for the nearest-neighbor model. For the long-range model, our RSRG-X simulation suffers from strong finite-size effects, but we observe the same asymptotic behavior at the largest accessible system size $N=40$ (Fig. 10). In both cases, our results suggest slower-than-power-law decay up to surprisingly large timescale in one dimension; recall that the typical energy scale (1) is fixed to unity. In two dimensions, we observe power-law, sub-diffusive decay of $\overline{S_{p}}(t)$ in the small system sizes to which we have access (Fig. 12). Further numerical and even analytical advances are necessary to determine the phenomenology in the thermodynamic limit.
@@
The presence of slow dynamics due to randomness has been explored in numerous theoretical works, including recent modifications of Lieb-Robinson bounds [56, 57] and proofs of sub-diffusive transport [58], both concerning disordered spin chains with general types of local interactions. One question to ask is whether the slower-than-power-law decay of spin survival probability we have observed in a family of spin chains with $\mathrm{U}(1)$ and $\mathbb{Z}_{2}$ symmetries suggests possibly tighter bounds on transport properties in this setting. Such constraints on transport, if any, might be advanced by our RSRG-X formalism combined with a careful analysis of the subleading contributions from the perturbative treatments; indeed the general idea that strong interactions renormalize in the first step to effectively weaker ones has been applied in Ref. [57]. Moreover, although Lieb-Robinson bounds have been derived for generic systems with long-range interactions [59, 60], our work raises questions over how these bounds are modified in systems with positional randomness.
@@
On a practical level, the wide range of interacting models to which our RSRG-X applies include systems with long-range random interactions of the form $XX+YY$ and also even $ZZ$. Many cutting-edge experimental platforms feature such interactions and thus can probe the signatures of spin dynamics predicted by our formalism. Conversely, an important application of our method is to calibrate and benchmark emerging hardware and algorithm for quantum computation. Especially, the quantum simulation of random one-dimensional spin models will be a valuable step forward: the dynamics in these models involve interference effects of a kind not seen in translation-invariant settings. Meanwhile, random systems in one dimension are typically not as intractable as in higher dimensions; even finding the ground state of a random classical Ising model is NP-complete in two dimensions [61]. The rapid progress in experimental capabilities offers the promise that theory and experiment on random quantum spin models can be more strongly coupled than they have been in the past, leading to new discoveries in this rich area of physics.
@@ Appendix B Interactions at two superspin-0 sites
In this appendix, we explain why we can expect the $X^{(0)}X^{(0)}$ and $Y^{(0)}Y^{(0)}$ types of interactions between two superspin-0 sites to be rarely the strongest in the effective Hamiltonian $H_{\mathrm{eff}}$ updated along RSRG-X. Our reasoning consists of two parts. First, we point out that the superspin-0 sites themselves are rare, so the strongest interaction rarely involves two such sites. Furthermore, we argue that the $X^{(0)}X^{(0)}$ and $Y^{(0)}Y^{(0)}$ interactions are weak if the three-site interactions, as we also expect, are weak. These theoretical considerations are further supported by the numerical RSRG-X of the specific models (1) we consider (Appendix E).
@@
To argue that the $X^{(0)}X^{(0)}$ and $Y^{(0)}Y^{(0)}$ interactions are weak as long as the three-site interactions in $H_{\mathrm{eff}}$ are weak, we consider how such interactions can be generated in some prior RG step. In fact, a superspin-0 site itself can only be generated when we branch into the $V_{-}^{\prime}$ subspace [case (ii) in Sec. V.1.3]: let the strongest interaction $H_{0}$ in $H_{\mathrm{eff}}$ at this step involve, without loss of generality, the sites $1$, $2$ of equal superspin-$\frac{m}{2}$ with $m>0$. Then, $H_{\mathrm{eff}}^{(V_{-}^{\prime})}$ contains in replacement of the sites 1, 2 a new superspin-0 site, which we label by 00, as well as new interactions of the form $X^{(0)}X^{(0)}$ and $Y^{(0)}Y^{(0)}$ — let us consider a new interaction, for example, $X_{0}^{(0)}X_{3}^{(0)}$. In light of the $\mathrm{U}(1)$ and $\mathbb{Z}_{2}$ symmetries (31), we exhaust the allowed interactions in $H_{\mathrm{eff}}$ among the sites 1, 2, 3,